# Outcomes truncated by death in RCTs: a simulation study on the survivor average causal effect

### v2

@@ 3.3 Mean squared error
The top part of Figure 3 shows the average MSE for each scenario and method with regard to $\theta_{1}$, with the 95 % Monte Carlo confidence interval. Likewise, the bottom part of Figure 3 shows the average MSE with regard to $\theta_{2}$. The average MSE only weakly depended on the treatment effect on the outcome, which can be seen from the very similar patterns in all three columns of Figure 3, but it strongly depended on the treatment effect on survival and on the method of analysis (rows and colors in Figure 3). As a consequence of the bias which was largest for complete case analysis, this method also had the largest average MSE of all methods in the presence of treatment effects on survival. The average MSE was similar and relatively small for all methods (and types of analyses) in scenarios without a treatment effect on survival (B, E, H), and largest in scenarios where Epo reduced survival compared to placebo (C, F, I, bottom rows of both parts of Figure 3). The latter is again a consequence of the larger mortality in these scenarios (Table 3), which also resulted in larger bias. As for bias, the average MSE of analyses using all covariates and those omitting head circumference was very similar, whereas the MSE of analyses omitting gestational age was increased, but still smaller than the MSE of complete case analysis.
@@ 3.4 Coverage
The top part of Figure 4 shows the coverage for each scenario and method with regard to $\theta_{1}$, with the 95 % Monte Carlo confidence interval. Likewise, the bottom part of Figure 4 shows the coverage with regard to $\theta_{2}$. Tang et al., (2005) suggested that coverage can be expected to lie within the interval of the nominal coverage probability $\pm$ standard error, which is the standard error of a proportion, i.e., $\sqrt{\frac{p(1-p)}{n}}$, here $\sqrt{\frac{0.95\cdot 0.05}{1300}}$, and that values below this interval can be interpreted as under-coverage and values above as over-coverage. In the context of our simulation study, under- and over-coverage mean that the corresponding estimand is captured less and more often, respectively. Accordingly, the acceptable coverage range is 0.938 to 0.962 in our case, which is shown in Figure 4 (dotted vertical lines). Similarly, we can check whether the nominal confidence-level of 0.95 is included in the observed 95 % Monte Carlo confidence intervals for coverage. Both conditions are met with respect to both estimands in most cases. Exceptions are (1) complete case analysis with regard to $\theta_{1}$ in scenario C and with regard to $\theta_{2}$ in scenario D, where the point estimate for coverage lies below the acceptable range and the 95 % CI for coverage lies almost entirely below 0.95, indicating under-coverage and (2) the SACE estimator using all covariates in scenario H with regard to $\theta_{1}$ and $\theta_{2}$, where the point estimate for coverage lies above the acceptable range and the 95 % CI for coverage lies entirely above 0.95, indicating over-coverage, and (3) for multiple imputation omitting head circumference in scenario H, similar to (2). Again as a consequence of the bias of complete case analysis, this method has a lower coverage than the other methods in the presence of treatment effects on survival. The coverage of the analyses with all covariates and those either omitting head circumference or gestational age did not show a very clear pattern, for example, coverage of analyses omitting gestational age was not always lower than coverage of the other analyses, except for scenarios with a negative treatment effect on the outcome.
@@ 4 Discussion
We investigated the setting of an RCT with a continuous outcome truncated by death and performed a simulation study to compare treatment effect estimates from complete case analysis, analysis after multiple imputation of the missing outcome values and the SACE estimator proposed by Hayden et al., (2005). Both alternative approaches to complete case analysis, although targeting different estimands (principal stratum estimand, hypothetical estimand), efficiently reduced the bias compared to complete case analysis and led to similar estimates of the treatment effect. While covariate information was used to reduce the analysis data set to the principal stratum of “always survivors” by the SACE estimator, it was was also used to substitute the outcome measurements truncated by death by multiple imputation. Further, we showed that results of the SACE estimator and multiple imputation were robust to minor violations of the “explainable nonrandom survival” or the “missing at random” assumption due to model misspecification (i.e., omission of the covariate head circumference at birth in the analysis although it was used to generate the data), but were less robust to stronger violations (i.e., omission of gestational age).
@@
Regarding the observability of principal strata, an interesting approach was used by Qu et al., (2023). They argued that the potential outcomes of patients under both treatments can be observed in a cross-over trial when carry-over and period effects can be excluded. Hence, they used a 2$\times$2 cross-over trial to evaluate commonly used assumptions to estimate principal stratum effects, including monotonicity, within-treatment principal ignorability, and cross-world assumptions of principal ignorability and principal strata independence. In line with our results in Table 4 in the supporting information, they found that the monotonicity assumption did not hold well for two principal stratum variables. Moreover, they found that “cross-world principal ignorability” and “cross-world principal stratum independence conditional on baseline covariates”, which they defined as two components of “explainable nonrandom survival” (as required by Hayden et al., , 2005), seemed reasonable.
@@
There are many other methods than the one suggested by Hayden et al., (2005) to estimate the SACE. We chose this method because it neither assumes monotonicity nor exclusion restriction, which we both found unrealistic for the EpoRepair trial. In fact, in the context of the SACE in an RCT, monotonicity may always be questionable because if one treatment is believed to benefit survival a priori, a clinical trial would be unethical (Wang et al., , 2017). However, avoiding monotonicity and exclusion restriction comes at the cost of making the “explainable nonrandom survival” assumption. As other cross-world assumptions, it is untestable using observed data, as it involves both potential outcomes of patients. Due to the simulation of both potential outcomes for each patient and subsequent ideal analysis using all covariates used in the simulation to model the survival probabilities, we created a situation in which the assumption was met. Further, we showed that results of the SACE estimator were robust to a minor violation of the the "explainable nonrandom survival" assumption due to omission of the covariate head circumference at birth in the analysis, but not to a stronger violation due to omission of the covariate gestational age. These results suggest that some degree of violation may be acceptable in practice. Omission of gestational age is quite an extreme scenario since it is probably the most obvious and important predictor of survival in preterm infants, and maybe to a slightly lesser degree, of cognitive outcome. Moreover, gestational age at birth is routinely available, at least in developed countries with sufficient monitoring during pregnancy. Omission of less important predictors (due to lack of knowledge or measurement) seems more realistic. Of note, the situation is not unlike that for methods to adjust for confounding in observational studies such as propensity score weighting, which require that no unexplained confounders exist to be unbiased, and this assumption is also untestable. In both situations, however, subject matter expertise can be used to ensure that the most important covariates/confounders are measured and used in the analysis. Further, the use of a cross-over trial to assess several assumptions of principal stratification methods by Qu et al., (2023) suggested that specifically the assumptions made by Hayden et al., (2005) may not be unreasonable. Other advantages of this specific SACE estimator are the actual estimation of the SACE together with a confidence interval and its relatively convenient applicability. Which assumptions are more or less realistic for a specific RCT should be assessed in each case. Further, it depends on the trial design and the available data (e.g., whether longitudinal data are available), which methods can be applied. Thus, although we used the SACE estimator of Hayden et al., (2005) in our simulation study, it is not our aim to specifically recommend this method. A whole plethora of methods for principal stratification can be found in Lipkovich et al., (2022) and references therein, some of which can be applied to estimate the SACE. For example, principal ignorability, wich is a less strong assumption than explainable nonrandom survival may be combined with the monotonicity assumption to estimate the SACE (Jo and Stuart, , 2009; Ding and Lu, , 2017; Bornkamp and Bermann, , 2019).
@@
Strengths of our simulation study are that we have written and made public a simulation study protocol in advance, and that we are comparing established methods, none of which we developed ourselves. A need for such neutral comparison studies has been identified, as studies comparing an own new method with existing methods are often overly optimistic with respect to the performance of the new method (Boulesteix et al., , 2013; Pawel et al., , 2024). A limitation of our simulation study is that we did not specifically assess situations in which the SACE ($\theta_{2}$) would differ from the causal effect of the treatment on the outcome in the absence of mortality ($\theta_{1}$). In fact, $\theta_{2}$ and $\theta_{1}$ were relatively similar in all scenarios, and as a consequence, the bias of the SACE estimator with regard to $\theta_{2}$ was only minimally smaller than the corresponding bias of the analysis using multiple imputation and vice versa. Simulation of scenarios in which these estimands differ may require a data generating mechanism that involves interactions between treatment and covariates regarding the outcome, when some of these covariates are also related to survival.

### v3

@@ 3.3 Mean squared error
The top part of Figure 3 shows the average MSE for each scenario and method with regard to $\theta_{1}$, with the 95 % Monte Carlo confidence interval. Likewise, the bottom part of Figure 3 shows the average MSE with regard to $\theta_{2}$. The average MSE only weakly depended on the treatment effect on the outcome, which can be seen from the very similar patterns in all three columns of Figure 3, but it strongly depended on the treatment effect on survival and on the method of analysis (rows and colors in Figure 3). As a consequence of the bias which was always largest for complete case analysis, this method also had the largest average MSE of all methods in the presence of treatment effects on survival. The average MSE was similar and relatively small for all methods (and types of analyses) in scenarios without a treatment effect on survival (B, E, H), and largest in scenarios where Epo reduced survival compared to placebo (C, F, I, bottom rows of both parts of Figure 3). The latter is again a consequence of the larger mortality in these scenarios (Table 3), which also resulted in larger bias with regard to $\theta_{1}$ and $\theta_{2}$. As for bias, the average MSE of analyses using all covariates and those omitting head circumference was very similar, whereas the MSE of analyses omitting gestational age was increased, but still smaller than the MSE of complete case analysis.
@@ 3.4 Coverage
The top part of Figure 4 shows the coverage for each scenario and method with regard to $\theta_{1}$, with the 95 % Monte Carlo confidence interval. Likewise, the bottom part of Figure 4 shows the coverage with regard to $\theta_{2}$. Tang et al., (2005) suggested that coverage can be expected to lie within the interval of the nominal coverage probability $\pm$ standard error, which is the standard error of a proportion, i.e., $\sqrt{\frac{p(1-p)}{n}}$, here $\sqrt{\frac{0.95\cdot 0.05}{1300}}$, and that values below this interval can be interpreted as under-coverage and values above as over-coverage. In the context of our simulation study, under- and over-coverage mean that the corresponding estimand is captured less and more often, respectively. Accordingly, the acceptable coverage range is 0.938 to 0.962 in our case, which is shown in Figure 4 (dotted vertical lines). Similarly, we can check whether the nominal confidence-level of 0.95 is included in the observed 95 % Monte Carlo confidence intervals for coverage. Both conditions are met with respect to both estimands in most cases. Exceptions are (1) complete case analysis with regard to $\theta_{1}$ in scenario C and with regard to $\theta_{2}$ in scenario D, where the point estimate for coverage lies below the acceptable range and the 95 % CI for coverage lies almost entirely below 0.95, indicating under-coverage and (2) the SACE estimator using all covariates in scenario H with regard to $\theta_{1}$ and $\theta_{2}$, where the point estimate for coverage lies above the acceptable range and the 95 % CI for coverage lies entirely above 0.95, indicating over-coverage, and (3) for multiple imputation omitting head circumference in scenario H, similar to (2). Again as a consequence of the bias of complete case analysis with regard to $\theta_{1}$ and $\theta_{2}$, this method has a lower coverage than the other methods in the presence of treatment effects on survival. The coverage of the analyses with all covariates and those either omitting head circumference or gestational age did not show a very clear pattern, for example, coverage of analyses omitting gestational age was not always lower than coverage of the other analyses, except for scenarios with a negative treatment effect on the outcome.
@@ 4 Discussion
We investigated the setting of an RCT with a continuous outcome truncated by death and performed a simulation study to compare treatment effect estimates from complete case analysis, analysis after multiple imputation of the missing outcome values and the SACE estimator proposed by Hayden et al., (2005). Both alternative approaches to complete case analysis, although targeting different estimands (principal stratum estimand, hypothetical estimand), efficiently reduced the bias compared to complete case analysis and led to similar estimates of the treatment effect. While covariate information was used to reduce the analysis data set to the principal stratum of “always survivors” by the SACE estimator, it was was also used to substitute the outcome measurements truncated by death by multiple imputation. Further, we showed that results of the SACE estimator and multiple imputation were robust to minor violations of the “explainable nonrandom survival” or the missing at random assumption due to model misspecification (i.e., omission of the covariate head circumference at birth in the analysis although it was used to generate the data), but were less robust to stronger violations (i.e., omission of gestational age).
@@
Regarding the observability of principal strata, an interesting approach was used by Qu et al., (2023). They argued that the potential outcomes of patients under both treatments can be observed in a cross-over trial when carry-over and period effects can be excluded. Hence, they used a 2$\times$2 cross-over trial to evaluate commonly used assumptions to estimate principal stratum effects, including monotonicity, within-treatment principal ignorability, and cross-world assumptions of principal ignorability and principal strata independence. In line with our results in Table 4 in the supporting information, they found that the monotonicity assumption did not hold well for two principal stratum variables. Moreover, they found that “cross-world principal ignorability” and “cross-world principal stratum independence conditional on baseline covariates”, which they defined as two components of “explainable nonrandom survival” (as required by Hayden et al., , 2005), seemed reasonable. However, this is evidence from a single study only, and it is contrasted by a recent theoretical assessment of these assumptions by Vansteelandt and Van Lancker, (2024), who judge them as implausible. In particular, assumption (7) is shown to be invalid under the null hypothesis of no treatment effect.
@@
There are many other methods than the one suggested by Hayden et al., (2005) to estimate the SACE. We chose this method because it neither assumes monotonicity nor exclusion restriction, which we both found unrealistic for the EpoRepair trial. In fact, in the context of the SACE in an RCT, monotonicity may always be questionable because if one treatment is believed to benefit survival a priori, a clinical trial would be unethical (Wang et al., , 2017). However, avoiding monotonicity and exclusion restriction comes at the cost of making the “explainable nonrandom survival” assumption. As other cross-world assumptions, it is untestable using observed data, as it involves both potential outcomes of patients. Due to the simulation of both potential outcomes for each patient and subsequent ideal analysis using all covariates used in the simulation to model the survival probabilities, we created a situation in which the assumption was met. Further, we showed that results of the SACE estimator were robust to a minor violation of the the "explainable nonrandom survival" assumption due to omission of the covariate head circumference at birth in the analysis, but not to a stronger violation due to omission of the covariate gestational age. These results suggest that some degree of violation may be acceptable in practice. Omission of gestational age is quite an extreme scenario since it is probably the most obvious and important predictor of survival in preterm infants, and maybe to a slightly lesser degree, of cognitive outcome. Moreover, gestational age at birth is routinely available, at least in developed countries with sufficient monitoring during pregnancy. Omission of less important predictors (due to lack of knowledge or measurement) seems more realistic. Of note, the situation is not unlike that for methods to adjust for confounding in observational studies such as propensity score weighting, which require that no unexplained confounders exist to be unbiased (e.g. with respect to the average treatment effect), and this assumption is also untestable. In both situations, however, subject matter expertise can be used to ensure that the most important covariates/confounders are measured and used in the analysis. Further, the use of a cross-over trial to assess several assumptions of principal stratification methods by Qu et al., (2023) suggested that specifically the assumptions made by Hayden et al., (2005) may not be unreasonable. Other advantages of this specific SACE estimator are the actual estimation of the SACE together with a confidence interval and its relatively convenient applicability. Which assumptions are more or less realistic for a specific RCT should be assessed in each case. Further, it depends on the trial design and the available data (e.g., whether longitudinal data are available), which methods can be applied. Thus, although we used the SACE estimator of Hayden et al., (2005) in our simulation study, it is not our aim to specifically recommend this method. A whole plethora of methods for principal stratification can be found in Lipkovich et al., (2022) and references therein, some of which can be applied to estimate the SACE. For example, principal ignorability, wich is a less strong assumption than explainable nonrandom survival may be combined with the monotonicity assumption to estimate the SACE (Jo and Stuart, , 2009; Ding and Lu, , 2017; Bornkamp and Bermann, , 2019).
@@
Strengths of our simulation study are that we have written and made public a simulation study protocol in advance, and that we are comparing established methods, none of which we developed ourselves. A need for such neutral comparison studies has been identified, as studies comparing an own new method with existing methods are often overly optimistic with respect to the performance of the new method (Boulesteix et al., , 2013; Pawel et al., , 2024). A limitation of our simulation study is that we did not specifically assess situations in which the SACE ($\theta_{2}$) would differ from the causal effect of the treatment on the outcome in the absence of mortality ($\theta_{1}$). In fact, $\theta_{2}$ and $\theta_{1}$ were relatively similar in all scenarios, and as a consequence, the bias of the SACE estimator with regard to $\theta_{2}$ was only minimally smaller than the corresponding bias of the analysis using multiple imputation and vice versa. Simulation of scenarios in which these estimands differ may require a data generating mechanism that involves interactions between treatment and covariates regarding the outcome, when some of these covariates are also related to survival. However, these results do not mean that the SACE and the hypothetical estimand are always similar, or that the methods to estimate one can be used to estimate the other.
@@
Acknowledgements
We thank two anonymous reviewers for their thorough reviews and very helpful comments.